OpenStudy (anonymous):

Determine the factored form of 7x^2+55x+42 using quadratic formula .

OpenStudy (anonymous):

x1=-6/7 and x2=-7

OpenStudy (anonymous):

So both of them are going to be negative ?

OpenStudy (anonymous):

quad formula x=[-b-+sqrtb^2-4ac]/2a

OpenStudy (anonymous):

yes both negative.try that formula

OpenStudy (anonymous):

Can you help on my other questions ?

OpenStudy (anonymous):

7x^2+55x+42 =0 x=[-b-+sqrt(b^2-4ac)]/2a x=[-55-+sqrt(55^2-4(7)(42))]/2(7) x=[-55-+sqrt(3025-1176))]/14 x=[-55-+sqrt(1849))]/14 x=[-55-+sqrt(1849))]/14 x=[-55-+43)]/14 x1=-98/14=-7 x2=[-55+43)]/14=-12/14=-6/7

OpenStudy (anonymous):

where is it